OpenStudy (anonymous):

radical 72x^5y^8 in simplest radical form

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

yes but x^5

OpenStudy (saifoo.khan):

\[\sqrt{72x^5y^8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh mistake

OpenStudy (anonymous):

72 = 9 * 4 * 2 radical 72 = 6 * 2^0.5 so : 6 * 2^0.5 * x^2.5 * y^4

OpenStudy (saifoo.khan):

c.green, are u doing this online? if yes, then this answer wont work.

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

coolsector, please write that in radical form.

OpenStudy (anonymous):

\[2\sqrt{6}x ^{2.5}y ^{4}\] that's what i get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh wait now i get what you need

OpenStudy (anonymous):

\[2\sqrt{6}\sqrt{x} x^2y^4\]